How do I find the answer to this problem. M=f/a for a. Also here's one more.
17 2/5=2 (3 1/2)+2L find the length.

Answer by Alan3354(69443)

How do I find the answer to this problem.
M=f/a for a. Also here's one more.
------------
Multiply by a.
Divide by M
=================================
17 2/5=2 (3 1/2)+2L find the length.
Not clear.